3g C reacts with 5g of O2 find percentage of CO in product?
 

Grade:11

1 Answers

Arun
5 years ago
The reaction for the same would be given by:
3C+2O2\rightarrow 2CO+CO2
Now the number of moles of C are equal to:
\frac{3}{12}=\frac{1}{4}
The number of moles of Oxygen are:
\frac{5}{32}
So the limiting reagent is Oxygen.
And so the mole fraction of CO2 would be equal to:
\frac{\frac{1}{2}}{\frac{1}{2}+\frac{1}{1}}=\frac{1}{3}
Regards
Arun